• Main INDEX
  • Monthly INDEX
  • PREV
  • NEXT

    User name Kramer

    Log entry time 23:56:10 on June21,2002

    Entry number 86553

    keyword=shift summary

    recovering from esr problems
    1600-1900 target recovery
    1900 start detector setup
    2000 problem with left dipole - current stuck, call in help
    2115 after cycling the left spectrometer power supply crate, gain control of supplies and start ramping
    2230 magnets back, but left dipole not up to full current. Stuck at 1265 A. Went to manual adjustment to get it to correct NMR. Final set point1369.42 for dipole. Q1 small adjustment to 2818.46 for setpoint.
    2245 took harp scan, raster off, 5uA. Analysis program not coming up or found. analyze later.
    2250 Run 1711 optics run
    2257 Move to D tgt, raster back on
    2306 start run 1712 production
    2334 Noticed that data acq rate dropped from 300 to 50 hz during run, even though beam, target, and scaler rates remain constant. Stop run 1712 and end of run failed. Restart coda.
    2337 start run 1713, coda back. 290 Hz